OpenStudy (superhelp101):

Use complete sentences to analyze the graph of the function f(x) = log 1/4 x. Part 1 : Describe the domain, range, and general shape of this graph. Part 2 : Using complete sentences, explain how plotting specific points helps graph the function and note any critical points such as its intercepts

OpenStudy (solomonzelman):

The relation is like this: \(\normalsize\color{blue}{ \log_{\large A}B=C~~~~~⇒~~~~~A^{\large C}=B }\)

OpenStudy (solomonzelman):

1/4 is the base right ?

OpenStudy (superhelp101):

yes

OpenStudy (solomonzelman):

So, then \(\normalsize\color{blue}{ y=\log_{\Large \frac{1}{4}}x~~~~~⇒~~~~~\Large{(}\Large \frac{1}{4})^{\large y}=x }\)

OpenStudy (superhelp101):

oh ok

OpenStudy (solomonzelman):

you can see that as x→∞, y→ 0 and when x→ -∞ y→ ∞

OpenStudy (superhelp101):

We learned the domains and ranges like for example: y<1.

OpenStudy (superhelp101):

wait, but I still not understanding part 2.

OpenStudy (superhelp101):

@SolomonZelman can you continue?

OpenStudy (superhelp101):

@paki @mathstudent55

OpenStudy (paki):

@SolomonZelman please have a look here...

OpenStudy (solomonzelman):

will just say that for example, (1/4) ^ y = x (1/4) ^ (1) = 1/4 and (1/4) ^ (2) = 1/16 and (1/4) ^ (3) = 1/64 This is what I was talking about when I said `y→∞, x→0`. The other way, (1/4) ^ (-1) = (4/1) ^ 1 = 4 (1/4) ^ (-2) = (4/1) ^ 2 = 16 (1/4) ^ (-3) = (4/1) ^ 3 = 64 and so on, saying that `y→ -∞, x→0` Y-intercept (1/4) ^ (0) = x x=1 X-intercept (1/4) ^ (y) = 0 y = no solution, because there is no number N, such that `(not a zero)^ N = zero.`